%merlin.mbs aipnum4-1.bst 2010-07-25 4.21a (PWD, AO, DPC) hacked
%Control: key (0)
%Control: author (8) initials jnrlst
%Control: editor formatted (1) identically to author
%Control: production of article title (0) allowed
%Control: page (1) range
%Control: year (1) truncated
%Control: production of eprint (0) enabled
\begin{thebibliography}{17}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{http://dx.doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\bibitem [{\citenamefont {Motta}\ \emph {et~al.}(2018)\citenamefont {Motta},
  \citenamefont {Ye}, \citenamefont {McClean}, \citenamefont {Li},
  \citenamefont {Minnich}, \citenamefont {Babbush},\ and\ \citenamefont
  {Chan}}]{motta2018low}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Motta}}, \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont {Ye}}, \bibinfo
  {author} {\bibfnamefont {J.~R.}\ \bibnamefont {McClean}}, \bibinfo {author}
  {\bibfnamefont {Z.}~\bibnamefont {Li}}, \bibinfo {author} {\bibfnamefont
  {A.~J.}\ \bibnamefont {Minnich}}, \bibinfo {author} {\bibfnamefont
  {R.}~\bibnamefont {Babbush}}, \ and\ \bibinfo {author} {\bibfnamefont
  {G.~K.}\ \bibnamefont {Chan}},\ }\bibfield  {title} {\enquote {\bibinfo
  {title} {Low rank representations for quantum simulation of electronic
  structure},}\ }\href@noop {} {\bibfield  {journal} {\bibinfo  {journal}
  {arXiv preprint arXiv:1808.02625}\ } (\bibinfo {year} {2018})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Whitten}(1973)}]{whitten1973coulombic}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.~L.}\ \bibnamefont
  {Whitten}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Coulombic
  potential energy integrals and approximations},}\ }\href@noop {} {\bibfield
  {journal} {\bibinfo  {journal} {The Journal of Chemical Physics}\ }\textbf
  {\bibinfo {volume} {58}},\ \bibinfo {pages} {4496--4501} (\bibinfo {year}
  {1973})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Chan}\ \emph {et~al.}(2016)\citenamefont {Chan},
  \citenamefont {Keselman}, \citenamefont {Nakatani}, \citenamefont {Li},\ and\
  \citenamefont {White}}]{chan2016matrix}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {G.~K.-L.}\
  \bibnamefont {Chan}}, \bibinfo {author} {\bibfnamefont {A.}~\bibnamefont
  {Keselman}}, \bibinfo {author} {\bibfnamefont {N.}~\bibnamefont {Nakatani}},
  \bibinfo {author} {\bibfnamefont {Z.}~\bibnamefont {Li}}, \ and\ \bibinfo
  {author} {\bibfnamefont {S.~R.}\ \bibnamefont {White}},\ }\bibfield  {title}
  {\enquote {\bibinfo {title} {Matrix product operators, matrix product states,
  and ab initio density matrix renormalization group algorithms},}\ }\href@noop
  {} {\bibfield  {journal} {\bibinfo  {journal} {The Journal of chemical
  physics}\ }\textbf {\bibinfo {volume} {145}},\ \bibinfo {pages} {014102}
  (\bibinfo {year} {2016})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Peng}\ and\ \citenamefont
  {Kowalski}(2017)}]{peng2017highly}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {B.}~\bibnamefont
  {Peng}}\ and\ \bibinfo {author} {\bibfnamefont {K.}~\bibnamefont
  {Kowalski}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Highly
  efficient and scalable compound decomposition of two-electron integral tensor
  and its application in coupled cluster calculations},}\ }\href@noop {}
  {\bibfield  {journal} {\bibinfo  {journal} {Journal of chemical theory and
  computation}\ }\textbf {\bibinfo {volume} {13}},\ \bibinfo {pages}
  {4179--4192} (\bibinfo {year} {2017})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Matsuzawa}\ and\ \citenamefont
  {Kurashige}(2020)}]{Matsuzawa2020}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont
  {Matsuzawa}}\ and\ \bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont
  {Kurashige}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Jastrow-type
  decomposition in quantum chemistry for low-depth quantum circuits},}\ }\href
  {\doibase 10.1021/acs.jctc.9b00963} {\bibfield  {journal} {\bibinfo
  {journal} {Journal of Chemical Theory and Computation}\ }\textbf {\bibinfo
  {volume} {16}},\ \bibinfo {pages} {944--952} (\bibinfo {year}
  {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Rubin}, \citenamefont {Lee},\ and\ \citenamefont
  {Babbush}(2021)}]{rubin2021compressing}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {N.~C.}\ \bibnamefont
  {Rubin}}, \bibinfo {author} {\bibfnamefont {J.}~\bibnamefont {Lee}}, \ and\
  \bibinfo {author} {\bibfnamefont {R.}~\bibnamefont {Babbush}},\ }\href@noop
  {} {\enquote {\bibinfo {title} {Compressing many-body fermion operators under
  unitary constraints},}\ } (\bibinfo {year} {2021}),\ \Eprint
  {http://arxiv.org/abs/2109.05010} {arXiv:2109.05010 [quant-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Helgaker}, \citenamefont {Jorgensen},\ and\
  \citenamefont {Olsen}(2014)}]{helgaker2014molecular}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {T.}~\bibnamefont
  {Helgaker}}, \bibinfo {author} {\bibfnamefont {P.}~\bibnamefont {Jorgensen}},
  \ and\ \bibinfo {author} {\bibfnamefont {J.}~\bibnamefont {Olsen}},\
  }\href@noop {} {\emph {\bibinfo {title} {Molecular electronic-structure
  theory}}}\ (\bibinfo  {publisher} {John Wiley \& Sons},\ \bibinfo {year}
  {2014})\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Landau}(1930)}]{Landau1930}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {L.}~\bibnamefont
  {Landau}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Diamagnetismus
  der metalle},}\ }\href {\doibase 10.1007/BF01397213} {\bibfield  {journal}
  {\bibinfo  {journal} {Zeitschrift f{\"u}r Physik}\ }\textbf {\bibinfo
  {volume} {64}},\ \bibinfo {pages} {629--637} (\bibinfo {year}
  {1930})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Fano}, \citenamefont {Ortolani},\ and\ \citenamefont
  {Colombo}(1986)}]{fano1986configuration}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {G.}~\bibnamefont
  {Fano}}, \bibinfo {author} {\bibfnamefont {F.}~\bibnamefont {Ortolani}}, \
  and\ \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont {Colombo}},\
  }\bibfield  {title} {\enquote {\bibinfo {title} {Configuration-interaction
  calculations on the fractional quantum hall effect},}\ }\href@noop {}
  {\bibfield  {journal} {\bibinfo  {journal} {Physical Review B}\ }\textbf
  {\bibinfo {volume} {34}},\ \bibinfo {pages} {2670} (\bibinfo {year}
  {1986})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Tsiper}(2002)}]{tsiper2002analytic}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Tsiper}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Analytic coulomb
  matrix elements in the lowest landau level in disk geometry},}\ }\href@noop
  {} {\bibfield  {journal} {\bibinfo  {journal} {Journal of Mathematical
  Physics}\ }\textbf {\bibinfo {volume} {43}},\ \bibinfo {pages} {1664--1667}
  (\bibinfo {year} {2002})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Kaicher}\ \emph {et~al.}(2020)\citenamefont
  {Kaicher}, \citenamefont {J{\"a}ger}, \citenamefont {Dallaire-Demers},\ and\
  \citenamefont {Wilhelm}}]{kaicher2020roadmap}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.~P.}\ \bibnamefont
  {Kaicher}}, \bibinfo {author} {\bibfnamefont {S.~B.}\ \bibnamefont
  {J{\"a}ger}}, \bibinfo {author} {\bibfnamefont {P.-L.}\ \bibnamefont
  {Dallaire-Demers}}, \ and\ \bibinfo {author} {\bibfnamefont {F.~K.}\
  \bibnamefont {Wilhelm}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Roadmap for quantum simulation of the fractional quantum hall effect},}\
  }\href@noop {} {\bibfield  {journal} {\bibinfo  {journal} {Physical Review
  A}\ }\textbf {\bibinfo {volume} {102}},\ \bibinfo {pages} {022607} (\bibinfo
  {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lloyd}(1996)}]{lloyd1996universal}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Lloyd}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Universal quantum
  simulators},}\ }\href {\doibase 10.1126/science.273.5278.1073} {\bibfield
  {journal} {\bibinfo  {journal} {Science}\ }\textbf {\bibinfo {volume}
  {273}},\ \bibinfo {pages} {1073--1078} (\bibinfo {year} {1996})},\ \Eprint
  {http://arxiv.org/abs/https://www.science.org/doi/pdf/10.1126/science.273.5278.1073}
  {https://www.science.org/doi/pdf/10.1126/science.273.5278.1073} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Jordan}\ and\ \citenamefont
  {Wigner}(1928)}]{Jordan1928}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {P.}~\bibnamefont
  {Jordan}}\ and\ \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Wigner}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {{\"U}ber das
  paulische {\"a}quivalenzverbot},}\ }\href {\doibase 10.1007/BF01331938}
  {\bibfield  {journal} {\bibinfo  {journal} {Zeitschrift f{\"u}r Physik}\
  }\textbf {\bibinfo {volume} {47}},\ \bibinfo {pages} {631--651} (\bibinfo
  {year} {1928})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Trotter}(1959)}]{trotter1959on}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.~F.}\ \bibnamefont
  {Trotter}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {On the product
  of semi-groups of operators},}\ }\href {http://www.jstor.org/stable/2033649}
  {\bibfield  {journal} {\bibinfo  {journal} {Proceedings of the American
  Mathematical Society}\ }\textbf {\bibinfo {volume} {10}},\ \bibinfo {pages}
  {545--551} (\bibinfo {year} {1959})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Wecker}\ \emph {et~al.}(2015)\citenamefont {Wecker},
  \citenamefont {Hastings}, \citenamefont {Wiebe}, \citenamefont {Clark},
  \citenamefont {Nayak},\ and\ \citenamefont {Troyer}}]{wecker2015solving}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {D.}~\bibnamefont
  {Wecker}}, \bibinfo {author} {\bibfnamefont {M.~B.}\ \bibnamefont
  {Hastings}}, \bibinfo {author} {\bibfnamefont {N.}~\bibnamefont {Wiebe}},
  \bibinfo {author} {\bibfnamefont {B.~K.}\ \bibnamefont {Clark}}, \bibinfo
  {author} {\bibfnamefont {C.}~\bibnamefont {Nayak}}, \ and\ \bibinfo {author}
  {\bibfnamefont {M.}~\bibnamefont {Troyer}},\ }\bibfield  {title} {\enquote
  {\bibinfo {title} {Solving strongly correlated electron models on a quantum
  computer},}\ }\href@noop {} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review A}\ }\textbf {\bibinfo {volume} {92}},\ \bibinfo {pages}
  {062318} (\bibinfo {year} {2015})}\BibitemShut {NoStop}%
\bibitem [{Note1()}]{Note1}%
  \BibitemOpen
  \bibinfo {note} {The leading order depends on the underlying qubit
  connectivity and whether one applies approximation techniques which further
  reduce the gate count, see Ref.~\cite {motta2018low}.}\BibitemShut {Stop}%
\bibitem [{\citenamefont {Schur}(1909)}]{Schur1909}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {I.}~\bibnamefont
  {Schur}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {{\"U}ber die
  charakteristischen wurzeln einer linearen substitution mit einer anwendung
  auf die theorie der integralgleichungen},}\ }\href {\doibase
  10.1007/BF01450045} {\bibfield  {journal} {\bibinfo  {journal} {Mathematische
  Annalen}\ }\textbf {\bibinfo {volume} {66}},\ \bibinfo {pages} {488--510}
  (\bibinfo {year} {1909})}\BibitemShut {NoStop}%
\end{thebibliography}%

\documentclass[%
 aip,
 amsmath,amssymb,
 reprint,%
]{revtex4-1}

\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{mathptmx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{nicefrac}
\usepackage{ulem}
\usepackage{qcircuit}
\usepackage{dcolumn}   
\usepackage{mathtools}
\usepackage{braket}
\usepackage{soul}
\usepackage{cancel}
\usepackage{mathdots}
\usepackage{graphicx}
\usepackage{placeins}
\usepackage{float}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage[bb=boondox]{mathalfa}
\usepackage{etoolbox}
\usepackage{dsfont}

\makeatletter
\def\@email#1#2{%
 \endgroup
 \patchcmd{\titleblock@produce}
  {\frontmatter@RRAPformat}
  {\frontmatter@RRAPformat{\produce@RRAP{*#1\href{mailto:#2}{#2}}}\frontmatter@RRAPformat}
  {}{}
}%
\makeatother
\begin{document}

\preprint{AIP/123-QED}

\title[Low-rank decomposition for quantum simulations with complex basis functions]{Low-rank decomposition for quantum simulations with complex basis functions}
\author{Michael P. Kaicher}
\affiliation{Departamento de F\'isica Te\'orica, Universidad Complutense, 28040 Madrid, Spain}

\date{\today}

\begin{abstract}
Low-rank decompositions to reduce the Coulomb operator to a pairwise form suitable for its quantum simulation are well-known in quantum chemistry, where the underlying basis functions are real-valued. We generalize the result of Ref.~\onlinecite{motta2018low} to \textit{complex} basis functions $\psi_p(\mathbf r)\in\mathds C$ by means of the Schur decomposition and decomposing matrices into their symmetric and anti-symmetric components. This allows the application of low-rank decomposition strategies to general basis sets. 
\end{abstract}

\maketitle

\section{Introduction}
When simulating a second quantized Hamiltonian of a fermionic system with  a quantum computer, recent works employed low-rank decomposition strategies known from classical simulation algorithms to express the interaction terms as sums of squares of one-body operators \cite{whitten1973coulombic,motta2018low,chan2016matrix,peng2017highly,Matsuzawa2020,rubin2021compressing}, which allows for a simulation in terms of fermionic Gaussian unitaries and Ising-type interactions \cite{motta2018low}. With the exception of Ref.~\onlinecite{rubin2021compressing}, all such decomposition strategies rely on symmetry properties of the two-body matrix elements in the second-quantized form which result from the the underlying single-particle basis functions being real-valued. This in particular applies to molecular electronic structure type Hamiltonians, which in general do not require complex basis functions \cite{helgaker2014molecular}. 

However, other quantum systems exist that are better described by complex basis functions, where the decomposition strategy of Ref.~\onlinecite{motta2018low} can no longer be applied. One prominent example are Landau-level wave functions \cite{Landau1930} used to represent the Hamiltonian describing the fractional quantum Hall effect \cite{fano1986configuration,tsiper2002analytic,kaicher2020roadmap}. Our work shows how one can employ the low-rank representation of the Hamiltonian terms describing the two-body interaction systems represented by complex-valued single-particle basis functions. 

This paper is structured as follows. In Section~\ref{motta} we review the low-rank strategy introduced by Motta \textit{et al.}, extend their result to complex-valued basis functions in Section~\ref{low_rank}, and conclude in Section~\ref{sum}. 
\section{Low-rank representation for real-valued basis functions\label{motta}}
Second quantized time-independent Hamiltonians that describe a non-relativistic system of $N_f$ interacting fermions are typically of the following form,
\begin{align}
\hat H=\hat H_1+\hat H_2,\label{mn1}    
\end{align}
where $\hat H_1$ ($\hat H_2$) contains a linear combination of a quadratic (quartic) polynomial of fermionic annihilation and creation operators, more specifically
\begin{align}
    \hat H_1 =& \sum_{p,q=1}^{N_f}f_{p,q} \hat c_p^\dag \hat c_q,\label{mn2}\\
    \hat H_2 =& \frac{1}{2}\sum_{p,q,r,s=1}^{N_f}h_{p,q,r,s} \hat c_p^\dag \hat c_q^\dag \hat c_r\hat c_s,\label{mn3}
\end{align}
where $f$ is a two-dimensional tensor describing the one-body terms such as kinetic energy or local potentials and $h$ is a four-dimensional tensor describing the physical (e.g. Coulomb-) interaction. Here, $\hat c_p^\dag$ and $\hat c_q$ are fermionic creation and annihilation operators acting on the spin-orbitals $p$ and $q$, $f_{p,q}$ and $h_{p,q,r,s}$ are overlap integrals where the subscript $p$ coressponds to the $p$-th single particle function $\psi_p(\mathbf r)$. The factor $\frac{1}{2}$ corrects for counting the interaction between two indistinguishable fermions twice. The tensors $f$ and $h$ carry a structure that depends on the employed basis functions, the quantum system at hand, and naturally reflect the fermionic nature of the problem. 

In this section, we will present a condensed version of the method described in Ref.~\onlinecite{motta2018low}, where $\hat H$ (using a real-valued basis set) can be rewritten as the sum-of-squares of one-particle operators, which allows for a relatively simple quantum simulation of its exponential map, e.g. required to simulate its time evolution \cite{lloyd1996universal}. 

For sets of \textit{real-valued} basis functions $\psi_p(\mathbf r)\in\mathds R$, explicit expressions were given for rewriting the Hamiltonian in Eq.~\eqref{mn1} as 
\begin{align}
\hat H =& \sum_{p,q} f_{p,q} \hat c_p^\dag \hat c_q+\sum_{p,q} S_{p,q} \hat c_p^\dag \hat c_q+\frac{1}{2} \sum_{L=1}^{N_f^2}\sum_{i,j}\lambda_i^{(L)}\lambda_j^{(L)}\hat n_i^{(L)}\hat n_j^{(L)}\nonumber\\
\equiv & \hat F +  \hat S + \sum_{L}\hat V^{(L)},\label{nm4}
\end{align}
where $f$ ($S$) is a two-dimensional tensor containing all one-body contributions (from the interaction term), $\hat n_i^{(L)}$ are number operators in a rotated basis, and $\lambda_i^{(L)}$ are coefficients. In fact, the coefficients $\lambda_i^{(L)}$ are connected to the spectrum of the reshaped tensor $h$ and can be used to truncate the summation over $L$, which can lead to a significant reduction in simulation cost at low truncation error \cite{motta2018low}.  Importantly, the creation and annihilation operators in the rotated basis only satisfy the anticommutation relations for a fixed $L$ and not for $L'\neq L$. This implies that $\left[\hat V^{(L)},\hat V^{(L')}\right]\neq 0$, in general. 

As one can see from Eq.~\eqref{nm4}, the simplicity of its quantum simulation becomes evident when applying the Jordan-Wigner transformation \cite{Jordan1928}. Then, the fermionic number operators (after applying the single-particle transformation $\hat U_L$) is diagonal with no appearance of Pauli-$Z$ strings, allowing for a simple simulation by means of an Ising-type interaction.

The exponential map of the Hamiltonian operator (e.g. for performing a unitary time evolution step) can than be approximated by a Trotter step  to first order \cite{trotter1959on}, 
\begin{align}
    e^{i\Delta t\hat H} \approx& e^{i\Delta t \left(\hat F + \hat S\right)}  \prod_{L=1}^{N_f^2}\left( \hat U^{(L)} e^{i\Delta t \hat V^{(L)}}\left(\hat U^{(L)}\right)^\dag\right),\label{mn5}
\end{align}
where $\hat U^{(L)}$ are basis rotation operations which can be implemented through Givens rotations \cite{wecker2015solving}. As discussed in Ref.~\onlinecite{motta2018low}, all terms on the right-hand side can be implemented using a low-order polynomial number of Givens rotations and phase gates \footnote{The leading order depends on the underlying qubit connectivity and whether one applies approximation techniques which further reduce the gate count, see Ref.~\cite{motta2018low}.}. The purpose of this work is to give the explicit expressions for the basis rotations for the case that the underlying single-particle basis functions in Eq.~\eqref{mn1} are complex-valued. 
\section{Factorization of the interaction term for complex-valued basis functions\label{low_rank}}
In the following, we will make extensive use of flattening (also known as \textit{reshaping}) of tensors, which is why we will be rigorous with comma-separated notation in the subscripts, e.g. $h_{p,q,r,s}$ is a $(N_f\times N_f\times N_f\times N_f)$-tensor, while $h_{pq,r,s}$ denotes the reshaped $(N_f^2\times N_f\times N_f)$-tensor, where the first two dimensions are flattened.  

 We consider the interaction term 
\begin{align}
   \hat H_2 = \frac{1}{2}\sum_{p,q,r,s=1}^{N_f}h_{p,q,r,s}\hat c_p^\dag \hat c_q^\dag \hat c_r \hat c_s, \label{p1059}
\end{align}
where the two-body matrix elements are defined as 
\begin{align}
h_{p,q,r,s}=&\frac{1}{4}\left( v_{p,q,r,s} - v_{q,p,r,s} + v_{q,p,s,r} -v_{p,q,s,r} \right),\label{int0}\\
v_{p,q,r,s}=&\iint d\mathbf r_1d\mathbf r_2\psi_p^*(\mathbf r_1)\psi_q^*(\mathbf r_2)\hat V(\mathbf r_1,\mathbf r_2)\nonumber\\&\times \psi_s(\mathbf r_1)\psi_r(\mathbf r_2).\label{int2}
\end{align}
Here, $\psi_p(\mathbf r)$ describes the $p$-th basis function for a particle located at position $\mathbf r$ and $V(\mathbf r_1,\mathbf r_2)$ describes the interaction potential between particle 1 and 2, typically given by the Coulomb potential $V(\mathbf r_1,\mathbf r_2)=1/|\mathbf r_1-\mathbf r_2|$, in appropriate units.  
We will assume that the resulting tensor elements computed through the two-electron integrals are real-valued, which leads to the property $h_{p,q,r,s}=(h_{p,q,r,s})^*=h_{s,r,q,p}$. One of the more prominent examples where this holds is for the fractional quantum Hall systems mentioned in the introduction, e.g. Haldane's spherical model \cite{fano1986configuration}, or the two-dimensional disk geometry \cite{tsiper2002analytic,kaicher2020roadmap}. Note, that since we have chosen $h_{p,q,r,s}$ to possess the following symmetries, $h_{p,q,r,s}=-h_{q,p,r,s}=-h_{p,q,s,r}=h_{q,p,s,r}$, this will also translate to $h_{s,r,q,p}$, resulting in the eight-fold symmetry
\begin{align}
    &h_{p,q,r,s}=-h_{q,p,r,s}=-h_{p,q,s,r}=h_{q,p,s,r}\nonumber\\ =& h_{s,r,q,p} = -h_{r,s,q,p} = -h_{s,r,p,q} = h_{r,s,p,q}.\label{p1078}
\end{align}
We rewrite Eq.~\eqref{p1059} as 
\begin{align}
   \hat H_2 = \frac{1}{2}\sum_{p,q,r,s=1}^{N_f}h_{p,q,r,s}\hat c_p^\dag  \hat c_s \hat c_q^\dag \hat c_r + \sum_{p,r=1}^{N_f}h_{p,r}\hat c_p^\dag  \hat c_r, \label{p1060}
\end{align}
where we defined 
\begin{align}
    h_{p,r}=-\frac{1}{2}\sum_{q=1}^{N_f}h_{p,q,r,q}.\label{p1061}
\end{align}
We begin by transposing the tensor, such that indices belonging to particle 1 (here $p,s$) and particle 2 (here $q,r$) are grouped together. This is followed by flattening the tensor into a $(N_f^2 \times N_f^2)$ matrix, so that $h_{p,q,r,s}=h_{ps,qr}$. Due to the symmetry properties in Eq.~\eqref{p1078}, we know that the flattened matrix $h_{ps,qr}$ is real and symmetric, which means that we can diagonalize it by means of a Schur decomposition \cite{Schur1909}, which results in
\begin{align}
    h = O^{[h]}\Sigma^{[h]}{O^{[h]}}^T,\label{p1079}
\end{align}
where $h$ is the flattened tensor with matrix elements $h_{ps,qr}$, $O^{[h]}$ is a $(N_f^2\times N_f^2)$-real orthogonal matrix, $\Sigma^{[h]}$ is a $(N_f^2\times N_f^2)$-diagonal matrix with non-negative real-valued entries, and the superscript $[h]$ indicates that $O$ and $\Sigma$ belong here to the decomposition of $h$. With this, we can write the first term on the right-hand side of Eq.~\eqref{p1060} as 
\begin{align}
\frac{1}{2}\sum_{ps,qr}^{N_f^2}h_{ps,qr}\hat c_p^\dag  \hat c_s \hat c_q^\dag \hat c_r
=&\frac{1}{2}\sum_{ps,qr}^{N_f^2}\sum_{L=1}^{N_f^2}\left(O^{[h]}\right)_{ps,L}\left(\Sigma^{[h]}\right)_{L,L}\nonumber\\&\times\left({O^{[h]}}\right)_{qr,L}\hat c_p^\dag  \hat c_s \hat c_q^\dag \hat c_r.\label{p1080}
\end{align}
We now introduce $L$-dependent matrices $O_L^{[h]}$ whose matrix elements are given by
\begin{align}
    \left(O_L^{[h]}\right)_{p,s} = \left(O^{[h]}\right)_{ps,L},\label{p1081}
\end{align}
which can easily be obtained from \texttt{numpy.reshape()}. We can then write Eq.~\eqref{p1080} as 
\begin{align}
&\frac{1}{2}\sum_{p,q,r,s=1}^{N_f}h_{p,q,r,s}\hat c_p^\dag  \hat c_s \hat c_q^\dag \hat c_r \nonumber\\
=&\frac{1}{2}\sum_{L=1}^{N_f^2}\sum_{ps,qr}^{N_f^2}\left(\Sigma^{[h]}\right)_{L,L}\left(O^{[h]}_L\right)_{p,s}\hat c_p^\dag  \hat c_s \left({O^{[h]}_L}\right)_{q,r}\hat c_q^\dag \hat c_r,\label{p1082}
\end{align}
and Eq.~\eqref{p1079} as 
\begin{align}
    h_{ps,qr} = \sum_{L=1}^{N_f^2}\Sigma^{[h]}_{L,L}\left(O^{[h]}_L\right)_{p,s}\left(O^{[h]}_L\right)_{q,r}.\label{p1083}
\end{align}
By introducing the real-valued symmetric and anti-symmetric components of the matrix $O_L^{[h]}$,
\begin{align}
    \mathcal S^{[O_L]}=&\frac{1}{2}\left(O^{[h]}_L+\left(O^{[h]}_L\right)^T\right),\label{p1084}\\
    \mathcal A^{[O_L]}=&\frac{1}{2}\left(O^{[h]}_L-\left(O^{[h]}_L\right)^T\right),\label{p1085}
\end{align}
we can write Eq.~\eqref{p1083} as 
\begin{align}
    h_{ps,qr} =& \sum_{L=1}^{N_f^2}\Sigma^{[h]}_{L,L}\left(\mathcal S^{[O_L]}\right)_{p,s}\left(\mathcal S^{[O_L]}\right)_{q,r}\nonumber\\&+\sum_{L=1}^{N_f^2}\Sigma^{[h]}_{L,L}\left(\mathcal A^{[O_L]}\right)_{p,s}\left(\mathcal A^{[O_L]}\right)_{q,r},\label{p1086}
\end{align}
leading to an expression for the interaction term in terms of symmetric and antisymmetric matrices. One might expect two additional terms appearing in Eq.~\eqref{p1086}, namely the cross terms $\mathcal S\mathcal A$ and $\mathcal A\mathcal S$, since Eq.~\eqref{p1086} is obtained by replacing $O^{[h]}_L$ in Eq.~\eqref{p1083} with its symmetric and anti-symmetric components defined in Eqs.~\eqref{p1084}-\eqref{p1085}. However, the cross terms vanish due to the symmetry constraint $h_{ps,qr}=h_{sp,rq}$, which follows from Eq.~\eqref{p1078}.  If we denote with $L_{\mathcal{S}}$ and $L_{\mathcal{A}}$ the set of indices in $L$ which give a non-zero symmetric matrix $\mathcal S^{[O_l]}$ and $\mathcal A^{[O_l]}$, respectively, with $L_{\mathcal{S}}\cup L_{\mathcal{A}}= \{1,2,\dots,N_f^2\}$ and $L_{\mathcal{S}}\cap L_{\mathcal{A}}=\{\}$, we can write Eq.~\eqref{p1086} as 
\begin{align}
    h_{ps,qr} =& \sum_{L_{\mathcal{S}}}\Sigma^{[h]}_{L_{\mathcal{S}},L_{\mathcal{S}}}\left(\mathcal S^{[O_{L_{\mathcal{S}}}]}\right)_{p,s}\left(\mathcal S^{[O_{L_{\mathcal{S}}}]}\right)_{q,r}\nonumber\\&+\sum_{L_{\mathcal{A}}}\Sigma^{[h]}_{L_{\mathcal{A}},L_{\mathcal{A}}}\left(\mathcal A^{[O_{L_{\mathcal{A}}}]}\right)_{p,s}\left(\mathcal A^{[O_{L_{\mathcal{A}}}]}\right)_{q,r},\label{p1087}
\end{align}
which leads to 
\begin{widetext}
\begin{align}
    \frac{1}{2}\sum_{p,q,r,s=1}^{N_f}h_{p,q,r,s}\hat c_p^\dag  \hat c_s \hat c_q^\dag \hat c_r 
=&\frac{1}{2}\sum_{L_{\mathcal{S}}}\sum_{p,q,r,s=1}^{N_f}\Sigma^{[h]}_{L_{\mathcal{S}},L_{\mathcal{S}}}\left(\mathcal S^{[O_{L_{\mathcal{S}}}]}\right)_{p,s}\hat c_p^\dag  \hat c_s \left(\mathcal S^{[O_{L_{\mathcal{S}}}]}\right)_{q,r}\hat c_q^\dag \hat c_r\nonumber\\
&+\frac{1}{2}\sum_{L_{\mathcal{A}}}\sum_{p,q,r,s=1}^{N_f}\Sigma^{[h]}_{L_{\mathcal{A}},L_{\mathcal{A}}}\left(\mathcal A^{[O_{L_{\mathcal{A}}}]}\right)_{p,s}\hat c_p^\dag  \hat c_s \left(\mathcal A^{[O_{L_{\mathcal{A}}}]}\right)_{q,r}\hat c_q^\dag \hat c_r. \label{p1088}
\end{align}
\end{widetext}
We have split the Hamiltonian into the symmetric and antisymmetric components of the sliced tensor of $O^{[h]}$ that diagonalizes the Coulomb tensor in its flattened form $h_{ps,qr}$ (Eq.~\eqref{p1079}). We can now diagonalize the symmetric matrix defined in Eq.~\eqref{p1084},
\begin{align}
     \mathcal S^{[O_L]}= U^{[\mathcal S^{[O_{L}]}]}\Sigma^{[\mathcal S^{[O_{L}]}]}{U^{[\mathcal S^{[O_{L}]}]}}^\dag,\label{p1089}
\end{align}
where $U^{[\mathcal S^{[O_L]}]}$ is a unitary matrix and $\Sigma^{[\mathcal S^{[O_L]}]}$ a real-valued diagonal matrix. Similarly, one can decompose the antisymmetric matrix defined in Eq.~\eqref{p1085} into
\begin{align}
    \mathcal A^{[O_L]}= U^{[\mathcal A^{[O_{L}]}]}\Sigma^{[\mathcal A^{[O_{L}]}]}{U^{[\mathcal A^{[O_{L}]}]}}^\dag,\label{p1090}
\end{align}
where $U^{[\mathcal A^{[O_L]}]}$ is a unitary matrix and $\Sigma^{[\mathcal A^{[O_L]}]}$ is a diagonal matrix only possessing purely imaginary entries. Note, that by definition $\mathcal S^{[O_{L_{\mathcal A}}]}=\mathbb 0_{N_f}=\mathcal A^{[O_{L_{\mathcal S}}]}$.

By introducing a new set of operators
\begin{align}
    \left.{\hat b}_a^{(L)}\right.^\dag =& \begin{cases}
    \sum_{p=1}^{N_f}\left(U^{[\mathcal S^{[O_{L_{\mathcal S}}]}]}\right)_{p,a}\hat c_p^\dag,\ \text{if}\ L\in L_{\mathcal S}, \\
    \sum_{p=1}^{N_f}\left(U^{[\mathcal A^{[O_{L_{\mathcal A}}]}]}\right)_{p,a}\hat c_p^\dag,\ \text{if}\ L\in L_{\mathcal A}, 
    \end{cases}\label{p1066}\\
    \hat b_a^{(L)} =& \begin{cases}
    \sum_{p=1}^{N_f}\left({U^{[\mathcal S^{[O_{L_{\mathcal S}}]}]}}^\dag\right)_{p,a}\hat c_p,\ \text{if}\ L\in L_{\mathcal S}, \\
    \sum_{p=1}^{N_f}\left(U^{[\mathcal A^{[O_{L_{\mathcal A}}]}]}\right)_{p,a}\hat c_p,\ \text{if}\ L\in L_{\mathcal A},
    \end{cases}\label{p1067}
\end{align}
we can express the Hamiltonian in terms of $L^2$ pairwise interaction terms. Note, that the operators defined in Eqs.~\eqref{p1066}-\eqref{p1067} only fulfill the canonical anticommutation relations for a fixed $L$, but not for $L'\neq L$. By defining $\hat n_a^{(L)}=\left.\hat b_a^{(L)}\right.^{\dag}\hat b_a^{(L)}$, we arrive at an expression of the interaction terms of Eq.~\eqref{nm4} for complex-valued basis functions, 
\begin{align}
    \sum_L\hat V^{(L)} =& 
    \frac{1}{2}\sum_{L_{\mathcal{S}}}\Sigma^{[h]}_{L_{\mathcal{S}},L_{\mathcal{S}}}\sum_{a,b=1}^{N_f} \Sigma^{[\mathcal S^{[O_L]}]}_{a,a}\Sigma^{[\mathcal S^{[O_L]}]}_{b,b}\hat n_a^{(L)}\hat n_b^{(L)}\nonumber\\
    &+\frac{1}{2}\sum_{L_{\mathcal{A}}}\Sigma^{[h]}_{L_{\mathcal{A}},L_{\mathcal{A}}}\sum_{a,b=1}^{N_f} \Sigma^{[\mathcal A^{[O_L]}]}_{a,a}\Sigma^{[\mathcal A^{[O_L]}]}_{b,b}\hat n_a^{(L)}\hat n_b^{(L)}.\label{nm10}
\end{align}
Inserting the right-hand side of Eq.~\eqref{nm10} into Eq.~\eqref{nm4} gives the low-rank representation of a second quantized Hamiltonian as in Eq.~\eqref{mn1} for complex-valued basis sets. This result was derived using only the fundamental anticommutation relation and indistinguishability of fermions, and assuming that the resulting overlap integrals are real-valued. 
\section{Conclusion\label{sum}}
In this work, we presented a low-rank decomposition of a general second quantized Hamiltonian with complex-valued basis functions into a sum of squared normal operators. For complex-valued basis functions, one loses a symmetry required to employ the method of Ref.~\onlinecite{motta2018low}, and we showed how this can be overcome by expressing the transformation matrices of the Schur-decomposed reshaped tensor $h$ in terms of its symmetric and antisymmetric components, and using the fermionic nature of the indistinguishable particles to cancel cross terms that mix anti-symmetric and symmetric components. This result allows one to apply low-rank decomposition-based quantum algorithms to general basis sets. Note, that at the time of this writing, a similar result was presented in Ref.~\cite{rubin2021compressing}, which also discusses how to determine a sum-of-squares decomposition of $\hat H_2$ by a greedy search algorithm, by means of a low-depth non-orthogonal one-particle bases expansion of $\hat H_2$. Since we did not perform numerical experiments and only mentioned truncation strategies to lower the number of elements (by e.g. truncating the sums over the symmetric and anti-symmetric components $L_{\mathcal S}$ and $L_{\mathcal A}$ by introducing a truncation threshold in the eigenvalues of $\Sigma$ as in Refs.~\onlinecite{motta2018low,rubin2021compressing}), future work should focus on explicit error analysis of such strategies to our decomposition. 
\begin{acknowledgments}
MPK thanks Simon Balthasar J\"ager for discussions. MPK acknowledges funding through S2018-TCS4342 QUITEMAD-CM.
\end{acknowledgments}

\nocite{*}
\bibliography{kaicher_JMP_submission_1}

\end{document}